Grade 11Physical Chemistry

45g of alloy is heated to a temperature of 90°C and dropped in 82g of water at temp 396.5 k.The temperature of water rises to 399.5 k ,given that the specific heat of water is 1Cal K-1 g-1 .Calculate the specific heat of alloy

Sabit
8 Years agoGrade 11

2 Answers

Khimraj
8 Years ago
As heat loss = heat gain
Khimraj
ApprovedApproved Tutor Answer8 Years ago
Q = mc(T2 – T1)
heat loss = heat gain
45*c*(399.5 – 363) = 82*1*(399.5 – 396.5)
So c = 0.15 Cal K-1g-1
Hope it clears.